# Rethinking Approximate Gaussian Inference in Classification

### v3

@@ 1 Introduction
For typical models $\bm{h}$ trained through such approximate maximum likelihood estimation, an output probability vector $\bm{h}(x)\in\Delta^{C-1}$ should be interpreted as an estimate of $p_{\text{gen}}(\bm{y}\mid x)$, the probability under the generative model. Notably, the output probability does not take into account the uncertainty of the model due to the finite nature of the data. In the uncertainty quantification formalism, such models can only estimate aleatoric uncertainty and disregard epistemic uncertainty [hullermeier2021aleatoric].
@@
The probabilistic way to capture epistemic uncertainty is to require the model to output a second-order distribution (that is ‘a distribution over probability distributions’). Namely, for each input $x\in\mathcal{X}$, the model should output a probability measure over the simplex $\mathrm{h}(x)\in\mathcal{P}(\Delta^{C-1})$, as opposed to a point estimate in the simplex $\bm{h}(x)\in\Delta^{C-1}$ [sensoy2018evidential, charpentier2020posterior].
@@
where $\mathrm{f}$ is learned from the training data, $\mathcal{G}(\mathbb{R}^{C})$ is the set of Gaussian measures on the logit space $\mathbb{R}^{C}$, and $\operatorname{\mathbf{softmax}}_{*}$ pushes forward the Gaussian probability measures through the softmax. For example, Heteroscedastic Classifiers (HET) [collier2021correlated] learn the output means and covariances of $\mathrm{f}$ explicitly with a neural network. In linearised Laplace approximations [daxberger2021laplace], only the means are learned explicitly with a neural network, while the covariances are obtained post-hoc from that neural network by approximating its parameter posterior distribution with a Gaussian, and then pushing it forward to logit space by linearising the network in the parameters. Last-layer Laplace approximations [kristiadi_being_2020] are a variant of linearised Laplace approximations where only the posterior over the last-layer parameters is Gaussian-approximated, requiring no linearisation of the network. Spectral-Normalised Gaussian processes (SNGP) [liu2020simple] are neural networks with last layers that are approximately Gaussian processes, through random feature expansions and Laplace approximations.
@@
In addition to predictive probabilities, the distributional framework allows for the acquisition of other quantities of interest, such as variances, entropy, and information-theoretic decompositions of the predictive uncertainty into aleatoric and epistemic parts [depeweg2018decomposition, mukhoti2023deep, wimmer2023quantifying]. These are usually obtained by Monte Carlo (MC) sampling from the Gaussian measure $\mathrm{f}(x)$. Indeed, while the probability density of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ can be obtained in closed form using a change of variables formula [atchison_logistic-normal_1980], this does not seem of practical use—even predictives cannot be computed analytically. On the other hand, given a variance, the runtime and memory cost required for an MC approximation to achieve an estimator of said variance grows linearly with the number of classes. Crucially, this computational cost is not limited to training time, but also to inference time, when the computational budget may be much more limited. This limits the application of such Gaussian inference methods in classification with a large number of classes, such as those common in computer vision and language modelling.
Summary of Contributions: Our work extends the classical probit approximation for the Gaussian sigmoid integral in binary classification [spiegelhalter_sequential_1990, mackay_evidence_1992] in the following manner:
@@
(i) Choose an approximate Gaussian inference method that outputs logit means and covariances. (ii) Choose an output activation that yields a closed-form Gaussian integral (Section˜3). (iii) At inference time, use our closed-form formulas (see Section˜3 for the predictives, Section˜4 for other quantities of interest).
@@ 2 Related Work
In uncertainty quantification, methods that output logit Gaussians, which we refer to as approximate Gaussian inference methods, have become one of the dominant paradigms in uncertainty quantification [collier2021correlated, daxberger2021laplace, kristiadi_being_2020, liu2020simple]. In classification tasks, such methods have the downside of requiring sampling at inference to transform Gaussian distributions into predictive probability vectors.
@@
In parallel, prior work has developed ways to propagate probabilistic uncertainty through neural networks [gast_lightweight_2018]. Several works develop analytic propagation of covariances up to the logits [wright_analytic_2024, li_streamlining_2024]; our method complements these by addressing the propagation of the logit-space Gaussian distributions onto the probability simplex. Few works address this last step. The Laplace bridge [pmlr-v180-hobbhahn22a] proposes a map from Gaussians to Dirichlets on the probability simplex. However, this approach has no guarantees on the quality of its predictives, and it severely underperforms Monte Carlo sampling. Other approaches include the mean-field approximation to the Gaussian softmax integral [lu_mean-field_2021] and the approximations of shekhovtsov_feed-forward_2018. In this work, we propose approximate predictives that outperform existing closed-form predictive techniques and provide second-order Dirichlet distributions for other closed-form quantities of interest. Furthermore, this is the first work to provide a theoretical analysis of such an approximation.
@@
Other works have also proposed various activations for classification to replace the softmax. Particularly relevant is [galy-fajou_multi-class_2020], which composes element-wise sigmoids with normalisation and utilises auxiliary-variable augmentation to obtain closed-form ELBO terms during training; however, its predictives require Monte Carlo estimation. Related conjugate/augmentation-based objectives exist for stick-breaking and tree-structured multi-class GPs [linderman_dependent_2015, achituve_gp-tree_2021], and for BNNs, a tempered-likelihood/Dirichlet perspective with a factorised-Gaussian surrogate has been explored [kapoor_uncertainty_2022]. Yet, all these approaches still require sampling or numerical quadrature at inference time. [spiegelhalter_sequential_1990, mackay_evidence_1992, kristiadi_being_2020] use the probit approximation in the binary case, which we generalise to the multi-class setting. [titsias_rc_aueb_one-vs-each_2016] approximates the softmax by providing lower bounds for point estimate training, but replacing the softmax with these bounds during training does not yield a normalised multi-class likelihood. The binary cross-entropy loss with the element-wise sigmoid activation has been reported to outperform cross-entropy in large-scale settings [wightman_resnet_2021, li_bce_2025]. Our work is the first to use this advantageous loss while normalising the output to obtain proper multi-class probabilistic models with closed-form predictives from logit Gaussians.
@@
Evidential deep learning (EDL) approaches [sensoy2018evidential] offer sample-free outputs through Dirichlet distributions but have been shown to be unreliable for uncertainty estimation and disentanglement [NEURIPS2022_bc1d640f, jürgens2024epistemicuncertaintyfaithfullyrepresented, pandey2023learn, mucsányi2024benchmarking]. We focus on making approximate Gaussian inference methods sample-free in a principled and performant manner; however, for completeness, we compare our approach with EDL methods in Section˜K.8.
@@
Finally, multiple authors have explored continuous approximations to categorical distributions, which enable gradient-based optimization for discrete stochastic models through sampling, such as the Gumbel-max trick [maddison_ast_2014, jang_categorical_2017, maddison_concrete_2017, huijben_review_2023]. These methods principally optimise a single categorical distribution, while the present work is concerned with constructing probability densities over the (continuous) simplex of such categorical distributions.
@@ 3.3 Theoretical Analysis of the Approximation’s Quality
$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})=M(\mathcal{K})+O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)$ (19)
as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$, for some constant $M(\mathcal{K})\in[0,\infty]$ which only depends on $\mathcal{K}$, not on $C$, such that $M(\mathcal{K})<\infty$ when $\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$, and $M(\mathcal{K})=0$ when $q=\hat{q}$ for all $(\mu,\sigma^{2})\in\mathcal{K}$.
Section˜C.2 provides a formula for $M(\mathcal{K})$ and a proof. We also conduct an empirical investigation of whether the assumptions made in the theorem hold in practice, namely whether the logit means and variances lie in a compact set, and whether $M(\mathcal{K})<\infty$.
The $O$ term in Eq.˜19 is due to the splitting of the expectation into the numerator and denominator in the derivation Eq.˜10, while $M(\mathcal{K})$ results from the error due to the probit approximation. Importantly, Theorem˜3.1 states that the latter does not increase with the number of classes.
@@ 4.1 Dirichlet Matching
where $T$ is Owen’s T function. For $\rho$, we have the following approximation (daunizeau_semi-analytical_2017)
@@
If we were to directly attempt to infer the parameters of the Dirichlet $\mathrm{h}(x)$ from Eqs.˜12 and 24, we would obtain an overparametrised system of equations with $2C$ equations and $C$ unknowns. Thus, we instead use a classical Dirichlet method of moments [minka_2000, Eqs. 19 & 23], which reduces these equations, as we will now describe.
@@ 5 Training
The BCE loss is a valid choice of loss as it is a (negative) strictly proper scoring rule (see Appendix˜I). Moreover, it has been observed to provide advantageous training dynamics compared to the CE loss [rw2019timm, li_bce_2025]. The choice of $\varphi=\Phi$ or $\rho$ not only fits our framework for closed-form predictives, but allows one to leverage such advantageous training dynamics (see experimental results in Section˜7).
@@ 7 Experiments
We consider Heteroscedastic Classifiers (HET) [collier2021correlated], Spectral-Normalised Gaussian Processes (SNGP) [liu2020simple], and last-layer Laplace-Approximated Networks [daxberger2021laplace] as backbones supplying logit-space Gaussian distributions. The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [deng2009imagenet], CIFAR-10, and CIFAR-100 [krizhevsky2009learning] on five metrics aligning with practical needs from uncertainty estimates [mucsanyi2023trustworthy]. Further, we test the Out-of-Distribution (OOD) detection capabilities of the models on balanced mixtures of In-Distribution (ID) and OOD samples. For ImageNet, we treat ImageNet-C [hendrycks2019benchmarking] samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10 and CIFAR-100, we use the CIFAR-10C corruptions.
For the second question, we evaluate our closed-form predictives and moment-matched Dirichlets against softmax models with the Laplace Bridge [pmlr-v180-hobbhahn22a] and Mean Field [lu_mean-field_2021] approximations, the linear-time approximation by shekhovtsov_feed-forward_2018, and MC sampling; see Section˜7.2.
To provide a fair comparison, we reimplement each method as simple-to-use wrappers around deterministic backbones. For ImageNet evaluation, we use a ResNet-50 backbone from the timm library [rw2019timm] pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following mucsányi2024benchmarking. For Vision Transformer [dosovitskiy2020image] experiments, refer to Section˜K.2. On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs. The CIFAR-100 experiments train WideResNet-28-5 models for 200 epochs. We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [shahriari2015taking] in Weights & Biases [wandb].
@@ 7.1 Quality of Sample-Free Predictives
In this section, we investigate whether there is a price to pay for sample-free predictives and second-order distributions. To this end, we take the two best-performing methods on the Expected Calibration Error (ECE) metric [naeini2015obtaining]—NormCDF and Sigmoid Laplace (see Fig.˜4(a))—and evaluate their per-predictive performance. As Table˜1 shows, the closed-form predictive is always on par with the (unbiased) MC estimates on ImageNet while being cheaper. Section˜K.5 reports the KL divergence of different approximations to the ‘true’ predictive (estimated using 10,000 MC samples) on ImageNet. Table˜2 shows that the moment-matched second-order Dirichlet distributions consistently outperform both MC and sample-free predictives on the OOD detection task, showcasing the practical utility of second-order distributional knowledge.
@@ 7.2 Effects of Changing the Learning Objective
We first evaluate how calibrated the models are using the test NLL [gneiting2007strictly] and Expected Calibration Error (ECE) [naeini2015obtaining] metrics. Fig.˜3(b) shows that on CIFAR-10, the NLL score of our closed-form predictives (Sigmoid, NormCDF) is consistently better than the corresponding Softmax results for all methods. On the large-scale ImageNet dataset, Sigmoid and NormCDF predictives either outperform or are on par with Softmax (see Fig.˜3(a)). The ECE metric requires the models’ confidence (i.e., maximum class probability) to match their accuracy. Fig.˜4(a) shows that on ImageNet, our closed-form predictives have a clear advantage over Softmax across all methods.
@@
Fig.˜5(a) evidences this claim on ImageNet: our closed-form predictives either outperform or are on par with Softmax predictives. These results are in line with the findings of wightman_resnet_2021, who also recommend training with a per-class binary cross-entropy loss using the sigmoid activation function. Section˜K.6 shows that the BCE loss already yields performance gains in the model backbones (without the logit-space Gaussians).
Finally, we consider another binary prediction task, where a general uncertainty estimator $u(x)\in\mathbb{R}$ (derived from predictives or second-order Dirichlet distributions) is tasked to separate ID and OOD samples from a balanced mixture thereof. As the uncertainty estimator can take on any real value, we measure the Area Under the Receiver Operating Characteristic curve (AUROC), which quantifies the separability of ID and OOD samples w.r.t. the uncertainty estimator. As OOD inputs, we consider corrupted ImageNet-C [hendrycks2019benchmarking] samples. Fig.˜5(b) shows that our closed-form predictives outperform Softmax across all methods. Importantly, these closed-form predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section˜6). See Section˜K.7 for the other severity levels.

### v4

@@ 1 Introduction
For typical models $\bm{h}$ trained through such approximate maximum likelihood estimation, an output probability vector $\bm{h}(x)\in\Delta^{C-1}$ should be interpreted as an estimate of $p_{\text{gen}}(\bm{y}\mid x)$, the probability under the generative model. Notably, the output probability does not take into account the uncertainty of the model due to the finite nature of the data. In the uncertainty quantification formalism, such models can only estimate aleatoric uncertainty and disregard epistemic uncertainty [21].
@@
The probabilistic way to capture epistemic uncertainty is to require the model to output a second-order distribution (that is ‘a distribution over probability distributions’). Namely, for each input $x\in\mathcal{X}$, the model should output a probability measure over the simplex $\mathrm{h}(x)\in\mathcal{P}(\Delta^{C-1})$, as opposed to a point estimate in the simplex $\bm{h}(x)\in\Delta^{C-1}$ [44, 6].
@@
where $\mathrm{f}$ is learned from the training data, $\mathcal{G}(\mathbb{R}^{C})$ is the set of Gaussian measures on the logit space $\mathbb{R}^{C}$, and $\operatorname{\mathbf{softmax}}_{*}$ pushes forward the Gaussian probability measures through the softmax. For example, Heteroscedastic Classifiers (HET) [7] learn the output means and covariances of $\mathrm{f}$ explicitly with a neural network. In linearised Laplace approximations [11], only the means are learned explicitly with a neural network, while the covariances are obtained post-hoc from that neural network by approximating its parameter posterior distribution with a Gaussian, and then pushing it forward to logit space by linearising the network in the parameters. Last-layer Laplace approximations [24] are a variant of linearised Laplace approximations where only the posterior over the last-layer parameters is Gaussian-approximated, requiring no linearisation of the network. Spectral-Normalised Gaussian processes (SNGP) [30] are neural networks with last layers that are approximately Gaussian processes, through random feature expansions and Laplace approximations.
@@
In addition to predictive probabilities, the distributional framework allows for the acquisition of other quantities of interest, such as variances, entropy, and information-theoretic decompositions of the predictive uncertainty into aleatoric and epistemic parts [13, 39, 53]. These are usually obtained by Monte Carlo (MC) sampling from the Gaussian measure $\mathrm{f}(x)$. Indeed, while the probability density of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ can be obtained in closed form using a change of variables formula [2], this does not seem of practical use—even predictives cannot be computed analytically. On the other hand, given a variance, the runtime and memory cost required for an MC approximation to achieve an estimator of said variance grows linearly with the number of classes. Crucially, this computational cost is not limited to training time, but also to inference time, when the computational budget may be much more limited. This limits the application of such Gaussian inference methods in classification with a large number of classes, such as those common in computer vision and language modelling.
Summary of Contributions: Our work extends the classical probit approximation for the Gaussian sigmoid integral in binary classification [47, 32] in the following manner:
@@
(i) Choose an approximate Gaussian inference method that outputs logit means and covariances. (ii) Choose an output activation that yields a closed-form Gaussian integral (Section˜3). (iii) At inference time, use our closed-form formulas (Section˜3 for the predictives, Section˜4 for other quantities of interest).
@@ 2 Related Work
In uncertainty quantification, methods that output logit Gaussians, which we refer to as approximate Gaussian inference methods, have become one of the dominant paradigms in uncertainty quantification [7, 11, 24, 30]. In classification tasks, such methods have the downside of requiring sampling at inference to transform Gaussian distributions into predictive probability vectors.
@@
In parallel, prior work has developed ways to propagate probabilistic uncertainty through neural networks [16]. Several works develop analytic propagation of covariances up to the logits [54, 28]; our method complements these by addressing the propagation of the logit-space Gaussian distributions onto the probability simplex. Few works address this last step. The Laplace bridge [19] proposes a map from Gaussians to Dirichlets on the probability simplex. However, this approach has no guarantees on the quality of its predictives, and it severely underperforms Monte Carlo sampling. Other approaches include the mean-field approximation to the Gaussian softmax integral [31] and the approximations of Shekhovtsov and Flach [46]. In this work, we propose approximate predictives that outperform existing closed-form predictive techniques and provide second-order Dirichlet distributions for other closed-form quantities of interest. Furthermore, this is the first work to provide a theoretical analysis of such an approximation.
@@
Other works have also proposed various activations for classification to replace the softmax. Particularly relevant is [15], which composes element-wise sigmoids with normalisation and utilises auxiliary-variable augmentation to obtain closed-form ELBO terms during training; however, its predictives require Monte Carlo estimation. Related conjugate/augmentation-based objectives exist for stick-breaking and tree-structured multi-class GPs [29, 1], and for BNNs, a tempered-likelihood/Dirichlet perspective with a factorised-Gaussian surrogate has been explored [23]. Yet, all these approaches still require sampling or numerical quadrature at inference time. [47, 32, 24] use the probit approximation in the binary case, which we generalise to the multi-class setting. [49] approximates the softmax by providing lower bounds for point estimate training, but replacing the softmax with these bounds during training does not yield a normalised multi-class likelihood. The binary cross-entropy loss with the element-wise sigmoid activation has been reported to outperform cross-entropy in large-scale settings [51, 27]. Our work is the first to use this advantageous loss while normalising the output to obtain proper multi-class probabilistic models with closed-form predictives from logit Gaussians.
@@
Evidential deep learning (EDL) approaches [44] offer sample-free outputs through Dirichlet distributions but have been shown to be unreliable for uncertainty estimation and disentanglement [4, jürgens2024epistemicuncertaintyfaithfullyrepresented, 42, mucsányi2024benchmarking]. We focus on making approximate Gaussian inference methods sample-free in a principled and performant manner; however, for completeness, we compare our approach with EDL methods in Section˜K.8.
@@
Finally, multiple authors have explored continuous approximations to categorical distributions, which enable gradient-based optimization for discrete stochastic models through sampling, such as the Gumbel-max trick [35, 22, 34, 20]. These methods principally optimise a single categorical distribution, while the present work is concerned with constructing probability densities over the (continuous) simplex of such categorical distributions.
@@ 3.3 Theoretical Analysis of the Approximation’s Quality
$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})\leq M(\mathcal{K})+O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)$ (19)
as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$ and $M(\mathcal{K})\in[0,\infty]$ some constant which only depends on $\mathcal{K}$, not on $C$, such that $M(\mathcal{K})<\infty$ when $\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$, and $M(\mathcal{K})=0$ when $q=\hat{q}$ for all $(\mu,\sigma^{2})\in\mathcal{K}$.
Section˜C.2 and Theorem˜C.2 provide a formula for $M(\mathcal{K})$ and a proof. We also conduct an empirical investigation of whether the assumptions made in the theorem hold in practice, namely whether the logit means and variances lie in a compact set, and whether $M(\mathcal{K})<\infty$.
The term $M(\mathcal{K})$ in Eq.˜19 results from the error due to the probit approximation, while the $O$ term is due to the splitting of the expectation into the numerator and denominator in the derivation Eq.˜10. Importantly, Theorem˜3.1 states that the former does not increase with the number of classes. For the latter, requiring $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$ is weaker than requiring $\mathrm{Var}(Q_{c})\to 0$ for all $c$ (see Eq.˜63 in Section˜C.2); but we show in Section˜C.2 that assuming $\mathrm{Var}(Q_{c})\to 0$ for all $c$ gives almost twice as strong a decay rate in $\operatorname{D_{KL}}$. We include this additional result in Theorem˜C.2.
@@ 4.1 Dirichlet Matching
where $T$ is Owen’s T function. For $\rho$, we have the following approximation (Daunizeau [10, Equation 23])
@@
If we were to directly attempt to infer the parameters of the Dirichlet $\mathrm{h}(x)$ from Eqs.˜12 and 24, we would obtain an overparametrised system of equations with $2C$ equations and $C$ unknowns. Thus, we instead use a classical Dirichlet method of moments [37, Eqs. 19 & 23], which reduces these equations, as we will now describe.
@@ 5 Training
The BCE loss is a valid choice of loss as it is a (negative) strictly proper scoring rule (see Appendix˜I). Moreover, it has been observed to provide advantageous training dynamics compared to the CE loss [52, 27]. The choice of $\varphi=\Phi$ or $\rho$ not only fits our framework for closed-form predictives, but allows one to leverage such advantageous training dynamics (see experimental results in Section˜7).
@@ 7 Experiments
We consider Heteroscedastic Classifiers (HET) [7], Spectral-Normalised Gaussian Processes (SNGP) [30], and last-layer Laplace-Approximated Networks [11] as backbones supplying logit-space Gaussian distributions. The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [12], CIFAR-10, and CIFAR-100 [25] on five metrics aligning with practical needs from uncertainty estimates [38]. Further, we test the Out-of-Distribution (OOD) detection capabilities of the models on balanced mixtures of In-Distribution (ID) and OOD samples. For ImageNet, we treat ImageNet-C [18] samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10 and CIFAR-100, we use the CIFAR-10C corruptions.
For the second question, we evaluate our closed-form predictives and moment-matched Dirichlets against softmax models with the Laplace Bridge [19] and Mean Field [31] approximations, the linear-time approximation by Shekhovtsov and Flach [46], and MC sampling; see Section˜7.2.
To provide a fair comparison, we reimplement each method as simple-to-use wrappers around deterministic backbones. For ImageNet evaluation, we use a ResNet-50 backbone from the timm library [52] pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following mucsányi2024benchmarking. For Vision Transformer [14] experiments, refer to Section˜K.2. On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs. The CIFAR-100 experiments train WideResNet-28-5 models for 200 epochs. We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [45] in Weights & Biases [5].
@@ 7.1 Quality of Sample-Free Predictives
In this section, we investigate whether there is a price to pay for sample-free predictives and second-order distributions. To this end, we take the two best-performing methods on the Expected Calibration Error (ECE) metric [40]—NormCDF and Sigmoid Laplace (see Fig.˜4(a))—and evaluate their per-predictive performance. As Table˜1 shows, the closed-form predictive is always on par with the (unbiased) MC estimates on ImageNet while being cheaper. Section˜K.5 reports the KL divergence of different approximations to the ‘true’ predictive (estimated using 10,000 MC samples) on ImageNet. Table˜2 shows that the moment-matched second-order Dirichlet distributions consistently outperform both MC and sample-free predictives on the OOD detection task, showcasing the practical utility of second-order distributional knowledge.
@@ 7.2 Effects of Changing the Learning Objective
We first evaluate how calibrated the models are using the test NLL [17] and Expected Calibration Error (ECE) [40] metrics. Fig.˜3(b) shows that on CIFAR-10, the NLL score of our closed-form predictives (Sigmoid, NormCDF) is consistently better than the corresponding Softmax results for all methods. On the large-scale ImageNet dataset, Sigmoid and NormCDF predictives either outperform or are on par with Softmax (see Fig.˜3(a)). The ECE metric requires the models’ confidence (i.e., maximum class probability) to match their accuracy. Fig.˜4(a) shows that on ImageNet, our closed-form predictives have a clear advantage over Softmax across all methods.
@@
Fig.˜5(a) evidences this claim on ImageNet: our closed-form predictives either outperform or are on par with Softmax predictives. These results are in line with the findings of Wightman et al. [51], who also recommend training with a per-class binary cross-entropy loss using the sigmoid activation function. Section˜K.6 shows that the BCE loss already yields performance gains in the model backbones (without the logit-space Gaussians).
Finally, we consider another binary prediction task, where a general uncertainty estimator $u(x)\in\mathbb{R}$ (derived from predictives or second-order Dirichlet distributions) is tasked to separate ID and OOD samples from a balanced mixture thereof. As the uncertainty estimator can take on any real value, we measure the Area Under the Receiver Operating Characteristic curve (AUROC), which quantifies the separability of ID and OOD samples w.r.t. the uncertainty estimator. As OOD inputs, we consider corrupted ImageNet-C [18] samples. Fig.˜5(b) shows that our closed-form predictives outperform Softmax across all methods. Importantly, these closed-form predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section˜6). See Section˜K.7 for the other severity levels.